OpenStudy (anonymous):

For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. x y 2 5 6 15 18 45

OpenStudy (anonymous):

For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. In the chart in the x slot its 2, 6 and 18. In the y its 5,15,and 45.

OpenStudy (tkhunny):

5/2 = 2.5 15/6 = 2.5 45/18 = 2.5 I'm convinced.

OpenStudy (anonymous):

Thanks :) Can you answer this one for me? For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. In the x slot is the numbers 16,32, and 48. In the y slot its 4,16 an 36.